OpenStudy (chosenmatt):

12+96 divided by 3 times 2 and 3 squared

OpenStudy (haichi):

72^3

OpenStudy (haichi):

12+96=108 and 108 divided by 3 is 36

OpenStudy (haichi):

36*2=72

OpenStudy (chosenmatt):

what i mean by squared is high 3 by 2

OpenStudy (haichi):

3 squared by 2 is 9

OpenStudy (chosenmatt):

so now i need help solving the prob

OpenStudy (haichi):

Ok

OpenStudy (haichi):

Use pemdas

OpenStudy (haichi):

3^2=9 9*3=27 108/27

OpenStudy (haichi):

108/27=4

OpenStudy (chosenmatt):

?

OpenStudy (haichi):

P=Parentheses E=Exponets M=Multiplication D=Division A=Add S=Subtract